# Towards Complex Ontology Alignment using Large Language Models

### v1

@@ 3.2 Design of the Prompting Process
Figure 2 illustrates our workflow, which begins by uploading the entire GMO file as an initial prompt. Subsequently, we include specific entities fom the second ontology, the GBO, to inquire about their alignment with the GMO. If the initial prompt successfully provides the relevant segments, further prompts in the chain-of-thought process are unnecessary. Otherwise, we supplement the inquiry with module information and request the related segments again.
@@ 4.1 Prompting and Detailed Example
To proceed, let us revisit the GMO modules discussed in Section 3.1. The documentation [14] contains an informal description of the modules (in the documentaiton called patterns) in the GMO ontology, accompanied by visual depictions of these patterns. In the ongoing chain-of-thoughts prompting processes, we included the module descriptions along with the question, as illustrated in the prompt below.
@@
By evaluating the outcome (see Fig. 6) of the prompt, it is evident that GPT-4 has a more informed and precise approach to conducting investigations and identifying semantically related components if it has the module information available. GPT-4 identifies all components related to the GMO as outlined in the alignment rule mentioned above. It further elaborates on each component and explains their interconnections. The full evaluation data of our study over all complex alignment rules is available online. [^8] https://tinyurl.com/geolinkComplexAlignmentEval
@@ 4.2 Quantitative Evaluation
detected pieces without module Information 2
detected pieces with module Information 107
@@
In Table 2 we see that for 67.6% of the complex alignment rules evaluated, the recall value exceeded 0.5. This indicates that more than half of the GMO entities involved in a complex alignment rule were successfully detected for this ratio in our population. Furthermore, the recall value surpasses 0.75 for approximately 57.4% of these rules, signifying a higher accuracy in detection. It’s also noteworthy that for 44.4% of the alignment rules recall is a perfect 1.0, and for 28.7% precision is also 1.0, because of the integration of module information in the analysis process.
The precision metric indicates the accuracy of the responses in directing us toward expected entities within the GMO ontology. Our analysis found that responses achieved a precision higher than 0.75 for 33% of the evaluated records. Additionally, when the precision threshold is lowered to 0.5—implying that half of the entities suggested by the language model are the expected ones, while the other half may include relevant or irrelevant entities—the coverage of alignment rules increases to 62%.
Note that detection (recall) of half or more of the correct body predicates, paired with high precision, is already immensely helpful for human assembly of a rule.
The recall data exhibits a mean of 0.68 and a median of 0.75, indicating that half of the recall values exceed 0.75, while the other half fall below this threshold. This distribution highlights that the majority of our data points demonstrate recall values above 0.75, offering insight into the central tendency of our dataset. Similarly, for precision, both the median and mean values suggest promising results, with at least 50% of the records achieving a precision greater than 0.5. This underscores a generally reliable level of accuracy in the data.
@@
with Module Information  67.5%   57.4%  44.4%  62%  33%  28.7%
mean 0.68 0.56
median 0.75 0.5
standard deviation 0.35 0.33

### v2

@@ 3.2 Design of the Prompting Process
Figure 2 illustrates our workflow, which begins by uploading the entire GMO file as an initial prompt. Subsequently, we include specific entities from the second ontology, the GBO, to inquire about their alignment with the GMO. If the initial prompt successfully provides the relevant segments, further prompts in the chain-of-thought process are unnecessary. Otherwise, we can provide a list of all GMO module names and ask GPT to identify the most related modules for the GBO entities. Then, in a subsequent prompt, we supplement the inquiry with module information and request the related segments again.
@@ 4.1 Prompting and Detailed Example
To proceed, let us revisit the GMO modules discussed in Section 3.1. The documentation [14] contains an informal description of the modules (in the documentation called patterns) in the GMO ontology, accompanied by visual depictions of these patterns. In the ongoing chain-of-thoughts prompting processes, we included the module names and asked for the most relevant module names related to the GBO pieces, as shown in Figure 6. The result of this prompt is usually a single module or a list of suggested modules based on the information that the GPT has processed through the chain of prompts. In the next prompt, we included the descriptions of the suggested modules along with the question, as illustrated in the prompt below.
@@
By evaluating the outcome (see Fig. 7) of the prompt, it is evident that GPT-4 has a more informed and precise approach to conducting investigations and identifying semantically related components if it has the module information available. GPT-4 identifies all components related to the GMO as outlined in the alignment rule mentioned above. It further elaborates on each component and explains their interconnections. The full evaluation data of our study over all complex alignment rules is available online. [^8] https://github.com/reddraa/complex_ontology_alignment
@@ 4.2 Quantitative Evaluation
detected pieces without module Information 5
detected pieces with module Information 104
@@
In Table 2 we see that for 73.3% of the complex alignment rules evaluated, the recall value exceeded 0.5. This indicates that more than half of the GMO entities involved in a complex alignment rule were successfully detected for this ratio in our population. Furthermore, the recall value surpasses 0.75 for approximately 62.3% of these rules, signifying a higher accuracy in detection. It’s also noteworthy that for 45% of the alignment rules recall is a perfect 1.0, and for 45.8% precision is also 1.0, because of the integration of module information in the analysis process.
The precision metric indicates the accuracy of the responses in directing us toward expected entities within the GMO ontology. Our analysis found that responses achieved a precision higher than 0.75 for 59.6% of the evaluated records. Additionally, when the precision threshold is lowered to 0.5—implying that half of the entities suggested by the language model are the expected ones, while the other half may include relevant or irrelevant entities—the coverage of alignment rules increases to 69.7%.
Note that detection (recall) of half or more of the correct body predicates, paired with high precision, is already very helpful for human assembly of a rule.
The recall data exhibits a mean of 0.67 and a median of 0.75, indicating that half of the recall values exceed 0.75, while the other half fall below this threshold. This distribution highlights that the majority of our data points demonstrate recall values above 0.75, offering insight into the central tendency of our dataset. Similarly, for precision, both the median and mean values suggest promising results, with at least 50% of the records achieving a precision greater than 0.8. This underscores a generally reliable level of accuracy in the data.
@@
with Module Information  73.3%   62.3%  45.0%  69.7%  59.6%  45.8%
mean 0.67 0.67
median 0.75 0.80
standard deviation 0.37 0.37